OpenStudy (anonymous):

How do you solve an indicated variable like 5y+2x=15 when y is the variable you need to solve for?

OpenStudy (amistre64):

algebra everything else to one side ...

OpenStudy (amistre64):

do you know how to solve an equation that has only 1 variable?

OpenStudy (anonymous):

Yes I do. But I was wondering if I'm supposed to divide the 5 at the same time I divide the 2x

OpenStudy (amistre64):

you cant really do two things at once here ... its just not that practical tell me why you think you need to divide a 2x?

OpenStudy (anonymous):

Well to solve for an indirect variable, you need to get that vaiable alone on one side. So if I was to get y on one side of the equal sign, then wouldn't I have to separate it from the 5 it's attached to? As well as the 2x? so the 5, 2x, and 15 would all some how equal y...

OpenStudy (amistre64):

if we work this step by step, not necessarily in this order ... but this is good enough 5y + 2x = 15 , to get rid the +2x, we subtract 2x from each side - 2x -2x ------------- 5y + 0x = 15 - 2x , simplfy the zero part 5y = 15 - 2x , now that we have a single term in y, we can now divide off that 5 /5 /5 ------------- 1y = (15-2x)/5 y = (15-2x)/5 , and simplfy to your hearts content

OpenStudy (anonymous):

oh wow, I COMPLETELY forgot that we had to do the opposite operation... so I was going to divide the 2x. Ugh first day of school review packet of something I did over 2 years ago without going over any of it isn't the best idea. why they put algebra 1, then geometry, then algebra 2 in that order, I'll never know. thanks for the help!

OpenStudy (amistre64):

:) good luck